If a 30-week deadline is imposed, what is the probability that the project will be finished within the time limit?


If the project manager wants to be 99% sure that the project is completed on the scheduled date, how many weeks before that date should he start the project work? (Please explain the answer of only this part of the question in brief)


Given Data:-

Variance of project (Sum of variance of critical activity) = 6

Expected project length of critical path = 5 + 15 + 4 + 5 = 29 weeks.



1
Expert's answer
2022-03-20T06:43:13-0400

"\mu=29\plusmn z_{\alpha/2}\frac{\sigma}{\sqrt{n}}=29 \plusmn 2.575 \frac{6}{\sqrt{4}}=29 \plusmn 7.725"

So, maximum it takes 37 weeks. If he wants make it for 30, he should start for 37-30=7 weeks before